OpenStudy (anonymous):

Calculate the velocity and acceleration vecdtors at t = 1 of r(t) = and then find the speed at t = 1

OpenStudy (anonymous):

Differentiate component-wise.$$\mathbf{r}'(t)=\left(2\cos2t,2\sin2t,\frac1{2\sqrt{t-1}}\right)\\\mathbf{r}''(t)=\left(-2\sin2t,2\cos2t,-\frac1{4\sqrt{(t-1)^3}}\right)$$Now evaluate at \(t=1\) for your velocity and acceleration vectors. Do you know how to determine the speed?

OpenStudy (anonymous):

oops$$\mathbf{r}''(t)=\left(-4\sin2t,4\cos2t,-\frac1{4(t-1)^\frac32}\right)$$

OpenStudy (anonymous):

To determine speed, isn't that the magnitude?

OpenStudy (anonymous):

of which vector?

OpenStudy (anonymous):

The velocity one!

OpenStudy (anonymous):

AKA 1st derivative

OpenStudy (anonymous):

Am I right @oldrin.bataku

OpenStudy (anonymous):

Indeed!